OpenStudy (priyar):

How to solve this problem? The velocity of water in a river is 18km/hr near the surface . If the river is 5m deep, find the shearing stress between the horizontal layers of water. The coeffecient of viscosity of water = 10^-2poise

OpenStudy (priyar):

@ParthKohli @imqwerty @ganeshie8

OpenStudy (priyar):

@UnkleRhaukus can u help?

OpenStudy (anonymous):

assuming Couette flow tau = mu *dv/dy I think

OpenStudy (priyar):

@ScientistsAreHot i didn't get u?

OpenStudy (baru):

\[\tau = \mu \frac{dv}{dy}\] is the formula @ScientistsAreHot is talking about

OpenStudy (baru):

the velocity at the bottom of the river is zero. you can assume that the velocity varies linearly from the bottom of the river to the top and write a function to give velocity at any height. the slope of this function is dv/dy. μ is viscosity.τ is shear stress.

OpenStudy (priyar):

i didn't know this formula...

OpenStudy (priyar):

how to write velocity as a function of ht @baru ?

OpenStudy (baru):



OpenStudy (baru):

V= (18/5) y

OpenStudy (baru):

sorry, you have to convert 18km/hr to m/s

OpenStudy (priyar):

ya i converted and multiplied by 0.1 since poise is 0.1kg(m)^-1(s)^-1 and i got the answer..! thanks a lot! can u help me with some more? i have tagged u..